# Relational Self-supervised Distillation with Compact Descriptors for Image Copy Detection

### v1

@@ Abstract.
This paper addresses image copy detection, a task in online sharing platforms for copyright protection. While previous approaches have performed exceptionally well, the large size of their networks and descriptors remains a significant disadvantage, complicating their practical application. In this paper, we propose a novel method that achieves a competitive performance by using a lightweight network and compact descriptors. By utilizing relational self-supervised distillation to transfer knowledge from a large network to a small network, we enable the training of lightweight networks with a small descriptor size. Our approach, which we call Relational self-supervised Distillation with Compact Descriptors (RDCD), introduces relational self-supervised distillation (RSD) for flexible representation in a smaller feature space and applies contrastive learning with a hard negative (HN) loss to prevent dimensional collapse. We demonstrate the effectiveness of our method using the DISC2021, Copydays, and NDEC benchmark datasets, with which our lightweight network with compact descriptors achieves a competitive performance. For the DISC2021 benchmark, ResNet-50/EfficientNet-B0 are used as a teacher and student respectively, the micro average precision improved by 5.0%/4.9%/5.9% for 64/128/256 descriptor sizes compared to the baseline method.
Knowledge Distillation, Self-supervised Learning, Lightweight Network, Image Copy Detection
1. Introduction
@@
Figure 1. Comparison of RDCD(Ours) and other image copy detection methods. RDCD utilizes a lightweight network and achieves high performance despite its compact descriptor sizes.
@@
Visual copy detection is a task that determines whether two pieces of multimedia items originate from a single source. This is widely used in online sharing platforms, including social networks, to filter content for copyright protection. Image copy detection (ICD), a part of visual copy detection, is the task of detecting copied images from an existing database, where each image instance is treated as an individual category such as instance matching problem (Douze et al., 2021a). This task has also been explored under the name of near-duplicate image detection (Chum et al., 2008) or content-based image retrieval (Kim, 2003). In this context, their approaches are centered on category-level recognition or instance-level recognition, focusing on identifying broad categories or specific objects, respectively. However, ICD addresses a more nuanced challenge by focusing on exact image copies that often involve severe transformations such as re-encoding, resizing, merging, cropping, warping, or color distortion. Traditional approaches in this domain have primarily employed either global descriptors (Kim, 2003; Wang and Jiang, 2015) or local descriptors such as bag-of-words (Zhou et al., 2010), which may struggle with such transformations. To better address these challenges, recent advancements have leveraged self-supervised learning (SSL) methods such as DINO (Caron et al., 2021), SimCLR (Chen et al., 2020b), MoCo (He et al., 2020), and Barlow-twins (Zbontar et al., 2021). These methods have been used to directly apply data augmentation as training objectives (Pizzi et al., 2022; Wang et al., 2021) to address severe transformations.
ICD involves two key challenges. The first is the large-scale detecting systems. In large-scale systems, millions of images in the database are pre-processed offline into descriptors and stored. Online query images are then converted into descriptors in real-time, and a nearest-neighbor search method is employed, similar to image search systems(Liu et al., 2007; Guo et al., 2020; Johnson et al., 2019). To address this issue, lightweight architectures have been introduced to convert images into descriptors in real-time. However, several studies (Shi et al., 2022; Gu et al., 2021) have found that lightweight networks with limited representation power cannot directly employ SSL to produce an acceptable performance. Previous studies have introduced SimCLR to train ResNet-50(RN-50) and ResNeXt-101, and we find that directly applying SimCLR to lightweight networks results in poor performance. Consequently, creating smaller descriptor sizes is an effective method to reduce the storage space for descriptors. Previous method (Pizzi et al., 2022) uses principal component analysis (PCA) to create small-sized descriptors, but it leads to significant performance degradation (Pizzi et al., 2022).
A second challenge for ICD stems from the difficulty of distinguishing hard negative samples. In ICD, there are numerous hard negative samples that are visually similar to the original image but are not edited copies. These include images taken from different camera angles or images of the same location captured at different times, which can significantly complicate the identification process. In ICD, each individual instance is treated as a separate category, making it crucial for the embedding space to achieve a uniform distribution. This arrangement helps to further each instance from its nearest neighbor (hard negative). In self-supervised copy detection (SSCD) (Pizzi et al., 2022), the Kozachenko-Leononenko estimator (Sablayrolles et al., 2018) is an entropy regularizer used to promote a uniform embedding distribution, thus ensuring that the distances from embedding regions are more comparable. By achieving a uniform distribution, the approach makes full use of all dimensions, thereby addressing the issue of dimensional collapse. Through experiments, we confirmed that training a student model with a teacher model previously trained using the entropy regularizer partially alleviated the issue of dimensional collapse. However, the student model still does not fully utilize all dimensions of the descriptors, continuing to struggle with dimensional collapse.
@@
In this work, we introduce Relational self-supervised Distillation with Compact Descriptors (RDCD), which utilizes both relation-based self-supervised distillation (RSD) and the HN loss. RSD is instrumental in accurately capturing the intricate relationships between teacher model descriptors, to ensure their effective representation in a reduced feature space. Existing SSL with knowledge distillation (SSL-KD) methods use feature-based knowledge distillation (FKD) (Gao et al., 2021) or RSD (Fang et al., [n. d.]; Abbasi Koohpayegani et al., 2020) for pre-training in classification tasks. FKD transfers knowledge by aligning embeddings of the student model with those of the teacher model using mean squared error (MSE) loss. However, we found that FKD fails to prevent dimensional collapse, leading to performance degradation in ICD. We create a separate instance queue for the teacher to make the student mimic the pairwise similarity generated by the teacher through Kullback-Leibler (KL) divergence, allowing the student network to generate discriminative descriptors in its own feature space.
To reduce the descriptor size, we use a linear projector to create descriptors and conduct SSL with MoCo-v2 (Chen et al., 2020a). It creates a multilayer perceptron (MLP) by adding two linear projectors: one to enable RSD with embeddings and the other for small descriptor sizes. According to (Jing et al., 2021), using an MLP in the last layer leads to dimensional collapse due to the multiplication of the weight matrices in each linear layer, which severely degrades overall performance. We use HN loss to avoid this collapse without impairing the performance of RSD from the teacher network, thus increasing the overall performance. The HN loss function is applied to the sample with the highest similarity within the instance queue used in MoCo-v2, to decrease the similarity of the most challenging negative samples. This function encourages the network to create more discriminative descriptors between positive and negative samples, leading to a more effective feature space where dissimilar samples are further apart. We consequently achieve a comparable performance using compact descriptors with our method as shown in Figure 1.
@@
We conduct comparative experiments with SimCLR (Chen et al., 2020b), MoCo-v2 (Chen et al., 2020a), FKD, and RSD to analyze the effectiveness of each method in training efficient lightweight networks for ICD. We test our approach on various teacher and student networks, including convolutional neural networks (CNNs) and vision transformers (ViTs), to demonstrate its architecture-agnostic characteristics. The main contributions of this work are summarized as follows:
@@
  • We introduce RDCD, a novel approach that combines RSD with HN loss to train lightweight networks for ICD in a self-supervised manner. Our method leverages the relational information between descriptors generated by a teacher network to guide the learning of a student network with a compact descriptor size.
  • Our RSD approach effectively maintains discriminative power when mapping to a reduced feature space, overcoming the limitations encountered in FKD.
@@
  • We employ HN loss to prevent the dimensional collapse that can occur in SSL with compact descriptors, ensuring a structured and informative feature space.
@@
2. Related Work
2.1. Knowledge Distillation
@@
Using a compact model to approximate a function learned by a more comprehensive, higher-performing model was first introduced in (Buciluǎ et al., 2006). This concept was expanded in (Hinton et al., 2015), where the student model was trained to mimic the teacher’s softened logits, a process referred to as knowledge distillation. Feature distillation has been garnering significant attention recently. (Romero et al., 2014) added a new dimension by proposing a hint-based training scheme for the alignment of feature maps, known as feature distillation, and chose the L2 distance as the metric for comparing the two feature maps. Similarly, (Zagoruyko and Komodakis, 2016) suggested the transfer of spatial attention maps from a high-performing teacher network to a smaller student network. (Yim et al., 2017) devised the idea of transferring ’flow’—defined as the inner product between features from two layers— to the student rather than knowledge. (Passalis and Tefas, 2018) directly aligned the probability distributions of the data between the teacher’s and student’s feature spaces. The success of these knowledge distillation methods can primarily be attributed to the insightful knowledge embedded in the logits of the teacher model.
2.2. Relation-based Knowledge Distillation
@@
While conventional knowledge distillation only extracts the information for a single data point from the teacher, similarity-based distillation methods (Tung and Mori, 2019; Abbasi Koohpayegani et al., 2020; Fang et al., [n. d.]; Passalis and Tefas, 2018; Park et al., 2019; Tian et al., 2019; Tejankar et al., 2021) learn the knowledge from the teacher in terms of similarities between data points. A pairwise similarity matrix has been proposed to retain the interrelationships of similar samples between the representation space of the teacher and student models. CompRess (Abbasi Koohpayegani et al., 2020) and SEED (Fang et al., [n. d.]) which are approaches that are most closely related to our method, use similarity-based distillation to compress a large self-supervised model into a smaller one.
2.3. Self-supervised Learning
In recent studies on SSL, contrastive-based approaches have produced remarkable results in downstream tasks. Many of these techniques are derived from noise-contrastive estimators (Gutmann and Hyvärinen, 2010), where the latent distribution is estimated based on the contrast with randomly or artificially generated noises. InfoNCE (Oord et al., 2018a) was the first to propose learning image representations by predicting the future using an auto-regressive model for unsupervised learning. Subsequent studies have focused on improving the efficiency (Henaff, 2020) and utilizing multi-view as positive samples (Tian et al., 2020). However, these approaches have limited access to negative instances, because they can only rely on a limited set of negative samples. To address this limitation, (Wu et al., 2018) designed a memory queue to store previously encountered random representations as negative samples, treating each of them as independent categories, for instance discrimination. However, this approach introduces an inefficiency in which the previously stored vectors are inconsistent with the recently computed representations during the earlier stages of pre-training. To mitigate this issue, SimCLR (Chen et al., 2020b) samples negative instances from a large batch while MoCo (He et al., 2020) improves on the memory-queue-based method by utilizing a momentum-updated encoder to reduce representation inconsistency.
Figure 2. Overall pipeline of proposed Relational Self-supervised Distillation for Image Copy Detection (RDCD).
3. Methodology
3.1. Overall Architecture
@@
The overall architecture of the proposed RDCD approach is presented in Fig 2. In RDCD, the student network is trained using three different objectives. Because the student network $f^{S}(\cdot)$ is the encoder that we want to improve, we freeze teacher network $f^{T}(\cdot)$ which is a pre-trained encoder by training using an off-the-shelf method.
@@
3.2. Relational Self-supervised Distillation
@@
The key objective is to align the similarity score distribution generated by the student network with that of the teacher network. This alignment is achieved by minimizing the KL divergence between the similarity score distributions of the two networks, ensuring a close match in their respective interpretations of the instance similarities. We apply the same augmentations into an image $x_{i}$ as a batch and map the embeddings $h^{T}=f^{T}_{\theta}(\tilde{x_{i}})$ and $h^{S}=f^{S}_{\theta}(\tilde{x_{i}})$ where $h^{T},h^{S}\in\mathbb{R}^{D}$ and $f^{T}_{\theta}$ and $f^{S}_{\theta}$ denote the teacher and student network, respectively.
@@
(1) $sim(h^{T}_{i},q^{T}_{j})=\begin{bmatrix}h^{T}_{i}/\lVert h^{T}_{i}\rVert_{2}\end{bmatrix}\cdot[q_{j}^{T}]^{\top}$
(2) $sim(h^{S}_{i},q^{T}_{j})=\begin{bmatrix}h^{S}_{i}/\lVert h^{S}_{i}\rVert_{2}\end{bmatrix}\cdot[q_{j}^{T}]^{\top}$
@@
(3) $p^{T}_{i,j}=\frac{exp(sim(h^{T}_{i},q^{T}_{j})/\tau^{T})}{\sum_{q^{T}\sim Q}exp(sim(h^{T}_{i},q^{T})/\tau^{T})},\quad where\;j\in[1\dots K]$
(4) $p^{S}_{i,j}=\frac{exp(sim(h^{S}_{i},q^{T}_{j})/\tau^{S})}{\sum_{q^{T}\sim Q}exp(sim(h^{S}_{i},q^{T})/\tau^{S})},\quad where\;j\in[1\dots K]$
@@
(5) $\qquad\log\left(\frac{\exp(\text{sim}(h^{s}_{i},q^{T}_{j})/\tau^{S})}{\sum_{q^{T}\sim Q^{T}}\exp(\text{sim}(h^{s}_{i},q^{T})/\tau^{S})}\right)$
3.3. Contrastive Learning for ICD
The size of the embeddings within the architecture of lightweight networks is generally small, and it is crucial to maintain this size during training while benefiting from knowledge distillation. To ensure this, we attach a linear projector to the end of the student network that is designed to reduce the representation size, resulting in smaller embeddings. We use these final embeddings as our descriptors, which allows for simultaneous training using our relational distillation approach. We follow the MoCo-v2 (Chen et al., 2020a) training procedure by utilizing a negative instance queue generated by the momentum encoder. We employ the InfoNCE (Oord et al., 2018b) loss to guide contrastive learning, which is calculated as follows:
(6) $\mathcal{L}_{con}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(z^{S}_{i},z^{S^{\prime}}_{i})/\tau)}{\sum_{q^{S}\sim Q^{S}}\exp(\text{sim}(z^{S}_{i},q^{S})/\tau)}$

### v5

@@ Abstract
Image copy detection is the task of detecting edited copies of any image within a reference database. While previous approaches have shown remarkable progress, the large size of their networks and descriptors remains a disadvantage, complicating their practical application. In this paper, we propose a novel method that achieves competitive performance by using a lightweight network and compact descriptors. By utilizing relational self-supervised distillation to transfer knowledge from a large network to a small network, we enable the training of lightweight networks with smaller descriptor sizes. We introduce relational self-supervised distillation for flexible representation in a smaller feature space and apply contrastive learning with a hard negative loss to prevent dimensional collapse. For the DISC2021 benchmark, ResNet-50 and EfficientNet-B0 are used as the teacher and student models, respectively, with micro average precision improving by 5.0%/4.9%/5.9% for 64/128/256 descriptor sizes compared to the baseline method. The code is available at https://github.com/juntae9926/RDCD.
@@
Figure 1: Comparison of RDCD(Ours) and other image copy detection methods. RDCD utilizes a lightweight network and achieves high performance with compact descriptor sizes.
1 Introduction
@@
Image copy detection (ICD) is the task of identifying edited copies of images within an existing database. This is widely used in online sharing platforms and social networks to filter content for copyright protection. Over the past years, ICD has been considered a sub-task of instance-level image retrieval [8]. These approaches utilize global descriptors [2, 33, 32, 24] and local descriptors [25, 37, 42], focusing on identifying specific objects that represent the images. However, ICD addresses a more complex challenge by focusing on detecting exact copies of the images themselves, rather than identifying representative objects. It deals with precise duplicates of images that undergo severe transformations such as re-encoding, resizing, merging, cropping, warping, or color distortion. To better address these challenges, recent studies have leveraged self-supervised learning (SSL) methods [5, 6, 15, 49] to directly apply data augmentation as their training objective [31, 45].
ICD involves two key challenges. The first challenge is the large scale of the detection system. In large-scale systems, millions of images in the database are pre-processed offline into descriptors and stored. Online query images are then converted into descriptors in real-time, after which a nearest-neighbor search method is employed, similar to image search systems [22, 14, 19]. However, this large-scale approach often leads to significant computational overhead, resulting in prolonged processing times, especially when dealing with vast image databases. To efficiently improve large-scale systems, an ideal approach is to use smaller architectures or utilize smaller-sized descriptors. This approach effectively reduces search time and storage space while maintaining system performance. However, several studies [36, 13] have found that lightweight networks with limited representation power cannot directly employ SSL to produce acceptable performance. Another method uses principal component analysis (PCA) to create small-sized descriptors, but it leads to significant performance degradation [31]. The second major challenge is the complex task of distinguishing between genuine copies and visually similar but unrelated images, known as hard negative samples. These hard negative samples are prevalent in ICD scenarios and can significantly complicate the identification process. Examples include images captured from different camera angles or images captured at the same location at different times, which are visually similar to the original but are not actual edited copies. A uniform distribution helps to maximize the separation between each image and its closest neighbors, including hard negatives. SSCD [31] employs the Kozachenko-Leononenko estimator [35] as an entropy regularizer to promote this uniform distribution. This technique ensures the distances between different regions in the embedding space are more consistent and comparable, thereby improving the system’s ability to differentiate between genuine copies and hard negative samples. By achieving uniform distribution, this approach makes full use of all dimensions, thereby addressing the issue of dimensional collapse. Despite these advancements, challenges still remain in efficiently training lightweight networks for ICD, especially when dealing with hard negatives, while maintaining high performance.
@@
In this work, we introduce Relational Self-supervised Distillation with Compact Descriptors (RDCD), a novel approach that combines relation-based self-supervised distillation (RSD) and Hard Negative (HN) loss to address these challenges. Our method aims to effectively capture the intricate relationships between teacher model descriptors and represent them in a reduced feature space. Existing SSL with knowledge distillation (SSL-KD) approaches use feature-based knowledge distillation (FKD) [12] or RSD [10, 1] for pre-training in classification tasks. FKD transfers knowledge by aligning embeddings of the student model with those of the teacher model using mean squared error loss. However, we found that FKD fails to prevent dimensional collapse, leading to performance degradation in ICD. To address this issue, we create a separate instance queue for the teacher to make the student mimic the pairwise similarity generated by the teacher using Kullback-Leibler (KL) divergence, allowing the student network to generate discriminative descriptors in its own feature space.
Although our distillation method is similar to previous SSL-KD methods [10, 1], it ultimately differs in the type of information it aims to distill. The main purpose of these methods is to provide advantages for downstream tasks as pre-training methods in lightweight networks. Their objective is to enrich the intermediate representations with transformation-covariant features, embedding useful information, such as color and orientation [6]. In contrast, our objective is to learn transformation-invariant representations at the final layer rather than obtaining transformation-covariant intermediate features. This design is crucial as the ICD task requires representations that go beyond instance discrimination to a tighter level, which is distinctly different from instance retrieval or classification tasks. In the ICD task, transformation-invariant features are essential, while intermediate representations with transformation-covariant characteristics degrade performance, as illustrated in [31].
@@
We conduct comparative experiments with SimCLR [6], MoCo-v2 [7], FKD, and RSD to analyze the effectiveness of each method in training efficient lightweight networks for ICD. We test our approach on various teacher and student networks, including convolutional neural networks and vision transformers, to demonstrate its architecture-agnostic characteristics. The main contributions of this work are summarized as follows:
@@
  • We introduce RDCD, a novel approach that combines RSD and HN loss to train lightweight networks for ICD in a self-supervised manner. Our method leverages the relational information between descriptors generated by a teacher network to guide the learning of a student network with a compact descriptor size.
@@
  • We employ HN loss to prevent the dimensional collapse that can occur in SSL using compact descriptors, ensuring a structured and informative feature space.
@@
Figure 2: The overall pipeline of proposed Relational Self-supervised Distillation for Image Copy Detection (RDCD). The method combines three key components: (1) Relational Self-supervised Distillation (RSD) which transfers knowledge from a pre-trained teacher network to a lightweight student network, (2) Contrastive Learning which uses MoCo-v2, and (3) Hard Negative (HN) Loss to address challenging negative samples. The student network $f^{S}_{\theta}$ extracts features $h^{S}$ and $h^{S^{\prime}}$ from two augmented views of an input image. These features are then projected to a lower dimension for use in contrastive learning. The teacher network $f^{T}_{\theta}$ guides RSD by comparing similarities between instances in its feature space. The final RDCD loss is a weighted combination of RSD loss ($\mathcal{L}_{rel}$), contrastive loss ($\mathcal{L}_{con}$), and HN loss ($\mathcal{L}_{hn}$), enabling the student to learn compact yet effective descriptors for image copy detection.
2 Related Work
2.1 Existing Image Copy Detection Methods
The ICD task has seen significant advancements since the 2021 Image Similarity Challenge [9] hosted by Meta AI Research, which spurred numerous research developments. Previous approaches utilize either contrastive self-supervised learning or deep metric learning. CNNCL [47] uses contrastive loss and cross-batch memory. EFNets [27] employs Arcface loss-based metric learning and uses a drip training technique to incrementally increase the number of classes during training. BOT [23] uses deep metric learning through hard sample mining with triplet loss and classification loss. In subsequent years, SSCD [31] applies InfoNCE loss with entropy regularization. Different from previous winning solutions which use numerous techniques to enhance performance, such as ensembles and post-processing, our approach follows the same evaluation dataset and basic evaluation methods from SSCD [31]. Additionally, we trained our model using a purely unlabeled dataset, without employing any ground truth labels. Active Image Indexing [11] applies imperceptible modifications to images called ”activation”, when using an ICD model in conjunction with Approximate Nearest Neighborhood indexing, ensuring the transformed image is positioned at the center of the quantized cluster. To address efficiency, our method aims to directly reduce descriptor sizes without relying on additional indexing techniques [38, 20].
2.2 Knowledge Distillation
@@
Using a compact model to approximate a function learned by a more comprehensive, higher-performing model was first introduced in [4]. This concept was expanded in [16], where the student model is trained to mimic the teacher’s softened logits, a process known as knowledge distillation. Feature distillation has been garnering significant attention recently. FitNet [34] adds a new dimension by proposing a hint-based training scheme for the alignment of feature maps, known as feature distillation, selecting the L2 distance as the metric for comparing the two feature maps. Similarly, [48] suggests the transfer of spatial attention maps from a high-performing teacher network to a smaller student network. [46] introduces the idea of transferring ’flow’—defined as the inner product between features from two layers— to the student. PKT [29] directly aligns the probability distributions of the data between the teacher’s and student’s feature spaces. The success of these knowledge distillation methods can primarily be attributed to the insightful knowledge embedded in the logits of the teacher model.
2.3 Relation-based Knowledge Distillation
@@
While conventional knowledge distillation only extracts the information for a single data point from the teacher, similarity-based distillation methods [43, 1, 10, 29, 28, 41, 39] learn the knowledge from the teacher in terms of similarities between data points. A pairwise similarity matrix has been proposed to retain the interrelationships of similar samples between the representation space of the teacher and student models. CompRess [1] and SEED [10], which are approaches related to our method, use similarity-based distillation to compress a large self-supervised model into a smaller one. However, our method differs in that we distill the relationship at the final layer to learn the transformation-invariant descriptor itself, whereas previous approaches enrich the intermediate representations with more transformation-covariant features.
3 Methodology
3.1 Overall Architecture
@@
The overall architecture of the proposed RDCD approach is presented in Fig 2. In RDCD, the student network is trained using three different objectives. Because the student network $f^{S}_{\theta}$ is the encoder that we want to improve, we freeze teacher network $f^{T}_{\theta}$ which is a pre-trained encoder by training using an off-the-shelf method.
@@
3.2 Relational Self-supervised Distillation
@@
The key objective is to align the similarity score distribution generated by the student network with that of the teacher network. This alignment is achieved by minimizing the KL-divergence between the similarity score distributions of the two networks, ensuring a close match in their respective interpretations of the instance similarities. We apply the same augmentations into an image $x_{i}$ as a batch and map the embeddings $h^{T}=f^{T}_{\theta}(\tilde{x_{i}})$ and $h^{S}=f^{S}_{\theta}(\tilde{x_{i}})$ where $h^{T},h^{S}\in\mathbb{R}^{D}$ and $f^{T}_{\theta}$ and $f^{S}_{\theta}$ denote the teacher and student network, respectively.
@@
$sim(h^{T}_{i},q^{T}_{j})=\begin{bmatrix}h^{T}_{i}/\lVert h^{T}_{i}\rVert_{2}\end{bmatrix}\cdot[q_{j}^{T}]^{\top}$ (1)
$sim(h^{S}_{i},q^{T}_{j})=\begin{bmatrix}h^{S}_{i}/\lVert h^{S}_{i}\rVert_{2}\end{bmatrix}\cdot[q_{j}^{T}]^{\top}$ (2)
@@
$p^{T}_{i,j}=\frac{exp(sim(h^{T}_{i},q^{T}_{j})/\tau^{T})}{\sum_{q^{T}\sim Q}exp(sim(h^{T}_{i},q^{T})/\tau^{T})},\;j\in[1\dots K]$ (3)
$p^{S}_{i,j}=\frac{exp(sim(h^{S}_{i},q^{T}_{j})/\tau^{S})}{\sum_{q^{T}\sim Q}exp(sim(h^{S}_{i},q^{T})/\tau^{S})},\;j\in[1\dots K]$ (4)
@@
$\qquad\log\left(\frac{\exp(\text{sim}(h^{s}_{i},q^{T}_{j})/\tau^{S})}{\sum_{q^{T}\sim Q^{T}}\exp(\text{sim}(h^{s}_{i},q^{T})/\tau^{S})}\right)$ (5)
3.3 Contrastive Learning for ICD
The size of the embeddings within the architecture of lightweight networks is generally small, and it is crucial to maintain this size during training while benefiting from knowledge distillation. To ensure this, we attach a linear projector to the end of the student network that is designed to reduce the representation size, resulting in smaller embeddings. We use these final embeddings as our descriptors, which allows for simultaneous training using our relational distillation approach. We follow the MoCo-v2 [7] training procedure by utilizing a negative instance queue generated by the momentum encoder. We employ the InfoNCE [26] loss to guide contrastive learning as follows:
$\mathcal{L}_{con}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(z^{S}_{i},z^{S^{\prime}}_{i})/\tau)}{\sum_{q^{S}\sim Q^{S}}\exp(\text{sim}(z^{S}_{i},q^{S})/\tau)}$ (6)